HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  laser spectroscopy on Be-like Kr  Not logged in ELOG logo
Message ID: 75     Entry time: Sun Mar 2 03:07:43 2025
Author: Denise, Benedikt, Danyal 
Category: DAQ 
Subject: 0.0012 nm shifted scan: 257.1738 nm - 257.6754 nm 
Zeit Scanbereich (nm) Anzahl Laserschritte Start LMD Stop LMD Messdauer (s) Elektronenkuehlung (s) Kommentar
02:47 Uhr 02.03.2025 257.1738 - 257.2986 52 krypton0195 krypton0202 10 30 -
03:38 Uhr 02.03.2025 257.2986 - 257.4234 52 krypton0203 krypton0210 10 30 -
04:31 Uhr 02.03.2025 257.4234 - 257.5482 52 krypton0211 krypton0218 10 30 -
05:25 Uhr 02.03.2025 257.5482 - 257.6754 53 krypton0219 krypton0226 10 30 -



Zeit Scanbereich (nm) Anzahl Laserschritte Start LMD Stop LMD Messdauer (s) Elektronenkuehlung (s) Kommentar
06:36 Uhr 02.03.2025 257.1738 - 257.2986 52 krypton0227 krypton0234 10 30 es gab ein paar leere Schüsse (keine Ionen)
07:29 Uhr 02.03.2025 257.1738 - 257.2986 52 krypton0235 krypton0242 10 30 voriges File wird deswegen wiederholt
08:46 Uhr 02.03.2025 257.2986 - 257.4234 52 krypton0243 krypton0250 10 30 -
09:41 Uhr 02.03.2025 257.4234 - 257.5482 52 krypton0251 krypton0257 10 30 -
10:34 Uhr 02.03.2025 257.5482 - 257.6754 53 krypton0258 krypton0265 10 30 -



Zeit Scanbereich (nm) Anzahl Laserschritte Start LMD Stop LMD Messdauer (s) Elektronenkuehlung (s) Kommentar
11:47 Uhr 02.03.2025 257.6754 - 257.8002 52 krypton0266 krypton0273 10 30
12:45 Uhr 02.03.2025 257.6754 - 257.8002 52 krypton0274 krypton0281 10 30
13:48 Uhr 02.03.2025 257.6754 - 257.8002 52 krypton0282 krypton0288 10 30


Zeit Scanbereich (nm) Anzahl Laserschritte Start LMD Stop LMD Messdauer (s) Elektronenkuehlung (s) Kommentar
14:50 Uhr 02.03.2025 257.5482 - 257.5962 20 krypton0289 krypton0291 10 30 scanned a small range where we thought there might be a peak in runs krypton0258 ff.


Zeit Scanbereich (nm) Anzahl Laserschritte Start LMD Stop LMD Messdauer (s) Elektronenkuehlung (s) Kommentar
15:28 Uhr 02.03.2025 257.0586 - 257.1738 48 krypton0292 krypton0297 10 30 there were three mis-injections from ESR in this scan
16:41 Uhr 02.03.2025 257.0586 - 257.1738 48 krypton0298 krypton0304 10 30 Signal drop at LaserStep = 34. I think I (Rodolfo) stopped the scan in LaserStep 47 (sorry)
17:34 Uhr 02.03.2025 257.0586 - 257.1738 48 krypton0305 krypton0311 10 30 there was a miss-injection in the ESR just at the end of the scan
ELOG V3.1.5-fc6679b